Circular No. 5885

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SPITALER (1993r) On Oct. 24 a comet was discovered by J. V. Scotti with the Spacewatch

telescope. He suggested that the object was P/Spitaler, not seen since its discovery in 1890, and for which S. Nakano (1992, Nakano Note No. 584) had predicted T = 1993 Oct. 4.0 TT. The comet was within a few arcmin of the line of variation at ΔT = +108.7 days. Follow-up observations were obtained on Oct. 26 and 27 by S. Larson with the 1.5-m Catalina reflector.

1993 UT α_{2000} δ_{2000} m_{1} Observer
Oct. 24.29677 1^{h}43^{m}06.^{s}34 +13^{o}35′43′′.3  Scotti
24.31248 1 43 05.54 +13 35 42.8 17.2  _{”}
24.42589 1 42 59.86 +13 35 37.2  _{”}
26.25626 1 41 32.66 +13 34 21.8  Larson
26.31402 1 41 29.76 +13 34 19.2  _{”}
27.32148 1 40 41.71 +13 33 34.9 17.5  _{”}
27.32421 1 40 41.57 +13 33 34.9  _{”}

J. V. Scotti (Kitt Peak). Nuclear mag m_{2} = 19.6. Coma diameter 18′′.

Tail 0′.76 long in p.a. 234^{o}. S. M. Larson (Lunar and Planetary Laboratory). CCD images measured

by C. Hergenrother, secondary references from the Oct. 24 CCD scans.

Computations by the undersigned confirm identity with P/Spitaler, perihelion dates for which are 1890 Oct. 27, 1897 Mar. 1, 1903 Dec. 15, 1910 Oct. 24, 1917 Aug. 30, 1924 July 8, 1931 May 21, 1938 Mar. 11, 1944 Dec. 18, 1951 Nov. 16, 1958 Nov. 18, 1965 Nov. 15, 1972 Nov. 6, 1979 Nov. 1, 1986 Dec. 26 and 1994 Jan. 28.

Epoch = 1994 Feb. 17.0 TT

T = 1994 Jan. 28.2251 TT  ω = 50.^{o}2047
e = 0.422305  Ω = 14.5105 }2000.0
q = 2.133171 AU  i = 5.7713
a = 3.692555 AU n^{o} = 0.1389037 P = 7.096 years
1993TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 20 1^{h}46.^{m}53 +13^{o}38.^{′}2 1.269 2.264 176.^{o}2 1.^{o}7 17.1
30 1 38.60 +13 31.6 1.255 2.240 170.6  4.2 17.0
Nov. 9 1 31.48 +13 24.6 1.265 2.218 159.1  9.2 17.0
19 1 26.32 +13 22.7 1.298 2.199 147.9 13.8 17.0
29 1 23.96 +13 30.5 1.351 2.182 137.4 17.8 17.0

1993 October 28 Brian G. Marsden